# Role of clustered nuclear geometry in particle production through p–C and p–O collisions at the Large Hadron Collider

### v1

@@ Abstract
Long-range multi-particle correlations in heavy-ion collisions have shown conclusive evidence of the hydrodynamic behavior of strongly interacting matter, and are associated with the final-state azimuthal momentum anisotropy. In small collision systems, azimuthal anisotropy can be influenced by the hadronization mechanism and residual jet-like correlations. Thus, one of the motives of the planned p–O and O–O collisions at the LHC and RHIC is to understand the origin of small system collectivity. As the anisotropic flow coefficients () are sensitive to the initial-state effects including nuclear shape, deformation, and charge density profiles, studies involving 12C and 16O nuclei are transpiring due to the presence of exotic (4He) clusters in such nuclei. In this study, for the first time, we investigate the effects of nuclear –clusters on the azimuthal anisotropy of the final-state hadrons in p–C and p–O collisions at TeV within a multi-phase transport model framework. We report the transverse momentum () and pseudorapidity () spectra, participant eccentricity () and triangularity (), and estimate the elliptic flow () and triangular flow () of the final-state hadrons using the two-particle cumulant method. These results are compared with a model-independent Sum of Gaussians (SOG) type nuclear density profile for 12C and 16O nuclei.
@@ I Introduction
The world’s most powerful particle accelerators, the Relativistic Heavy-Ion Collider (RHIC) at BNL, USA, and the Large Hadron Collider (LHC) at CERN, Switzerland, offer unparalleled experimental capabilities to perform high-energy hadronic and nuclear collisions to study the behavior of nuclear matter under extreme conditions of temperature and baryon density. Such ultra-relativistic heavy-ion collisions produce a state of locally thermalized and deconfined partonic medium, known as the quark-gluon plasma (QGP). While the existence of QGP has long been established in heavy-ion collisions, its presence in small collision systems is a matter of intense research, recently. Thanks to the recent measurements of heavy-ion-like features in high multiplicity pp collisions such as strangeness enhancement [1], ridge-like structure [3, 2], and radial flow effects [4, 5, 6], the possible existence of QGP droplets in small collision systems needs to be re-examined. As p–C and p–O collisions fill the multiplicity gap between pp to p–Pb and peripheral Pb–Pb collisions, it provides a perfect system size for studying several heavy-ion-like effects in small collision systems at the RHIC and LHC energies [7, 8].
@@
The anisotropic flow coefficients, such as elliptic flow () and triangular flow (), are found to be sensitive to the initial nuclear distribution, nuclear shape deformation, and fluctuations in the nuclear overlap region [16, 14, 15]. In Xe–Xe collisions at the LHC, the presence of a deformed nuclear structure gives rise to a higher value of elliptic flow in the most central collisions compared to Pb–Pb collisions [17, 18, 20, 19]. In addition, a larger density fluctuation due to a smaller number of participating nucleons in Xe–Xe collisions gives rise to a larger value of triangular flow in contrast to Pb–Pb collisions [17, 18]. A similar study using AMPT in U–U collisions at the RHIC energy shows that the values of anisotropic flow coefficients change with a change in the deformation parameter of the nuclear distribution [14]. However, the impact of this deformation is larger for elliptic flow in contrast to the triangular flow [14, 15, 21]. In addition, recent observations of ridge-like structures in p–Pb collisions have brought significant attention to asymmetric collisions, which are customarily considered to provide a baseline measurement for the cold nuclear effects [22, 23, 24, 25, 6], where the presence of strong long-range multi-particle correlations are not expected. Further, the measured values of anisotropic flow coefficients in p–Pb collisions are comparable to that of peripheral Pb–Pb collisions [27, 26, 28]. Although, in p–Pb collisions, the ordering of the magnitude of flow coefficients, i.e., , is not as strong as in Pb–Pb collisions, the bare presence of signatures of collectivity in p–Pb collisions makes the study of asymmetric collisions a vital contribution of the heavy-ion physics community.
@@
With all these tempting motivations, in recent years, several studies have been performed involving collisions of 12C and 16O nuclei [40, 41, 39, 35, 36, 37, 16, 38]. It includes the studies based on different hydrodynamic models [42, 43, 44], Glauber Monte Carlo [35, 36, 37], parton energy loss [45], and jet quenching effects [46]. Some of these studies also investigate the possible signatures of –clusters in 12C and 16O nuclei by studying the final-state particle production and anisotropic flow measurements [40, 41, 39, 48, 49, 50, 51, 16, 47]. In Ref. [40], an enhanced value of triangular flow in C–Au collisions is observed for an –clustered 12C nucleus. In Ref. [16], authors show that the presence of an –clustered structure in 16O nucleus leads to a high value of triangular flow in most central O–O collisions, which are well observed in the ratio . In addition, an away-side broadening in the two-particle correlation function is also reported for O–O collisions with –clusters as compared to the Woods-Saxon nuclear profile [16]. Although there have been a few studies that aim to establish a clear signature of the presence of –clusters in 16O and 12C nucleus, the choice of p–O and p–C is novel to this study. In addition, the studies of anisotropic flow in asymmetric collision systems like p–O and p–C can accord the observations of collectivity in p–Pb collisions.
@@
In this study, we incorporate an –cluster-type initial geometry for 12C and 16O nuclei, and report the first measurements of transverse momentum () and pseudorapidity () spectra, participant eccentricity () and triangularity (), and elliptic flow () and triangular flow () of the final-state hadrons, and their scaling in p–O and p–C collisions at TeV using a multi-phase transport model (AMPT). We use the two-particle cumulant method with a relative pseudorapidity cut to minimize the effects of nonflow correlations. These results are compared with a model-independent Sum of Gaussians (SOG) type nuclear density profile for 12C and 16O nuclei.
@@
The paper is organized as follows. We start with a brief introduction and motivation of the study in Section I followed by a description of event generation and methodology in Section II. In Section III, we present and discuss the results. Finally, in Section IV, we summarize the study with important findings and present a brief outlook for future research.
II Event generation and methodology
In this section, we begin with a brief introduction to a multi-phase transport model, which is used to simulate p–C and p–O collisions at TeV. The technical details of the implementation of different nuclear profiles for 16O and 12C nuclei are discussed next. Finally, a detailed description of the two-particle cumulant method for the anisotropic flow estimation is provided.
II.1 A multi-phase transport model
Since the impact parameter (), the number of participants (), and the number of binary collisions () cannot be directly measured in experiments, we resort to Glauber model estimations for the same [68, 69]. Using the publicly available MC Glauber code (TGlauberMC) [69], the impact parameter distribution is sliced to obtain the required centralities of collision. We have also modified the in-built HIJING model in AMPT to incorporate SOG and cluster nuclear density profiles for both and nuclei. Due to the unavailability of experimental data for p–O and p–C collisions at the LHC, we tune the AMPT model parameters by comparing them with that of the p–Pb system. A detailed description of the parameter tuning is provided in the Appendix.
II.2 Nuclear density profiles
@@
This study implements two different nuclear density profiles for the 12C and 16O nuclei, which are –cluster type geometrical distribution and a model-independent Sum of Gaussians type nuclear density profile. Here, the technical details regarding the implementation of the two density profiles are discussed.
II.2.1 Sum Of Gaussians (SOG) density profile
@@
To represent the nuclear charge density and extract charge density parameters in a model-independent fashion as approximated in experiments, one fits the nuclear densities with a sum of Gaussian functions (SOG) [70, 71]. SOG can accurately represent complex nuclear densities by exploiting the parameters of individual Gaussian functions, which makes SOG applicable to a broad list of nuclei. SOG can produce smooth and continuous descriptions of the nuclear density profiles, consequently simplifying many theoretical calculations due to the well-known properties of the Gaussian distribution function. In addition, SOG can produce a more realistic and precise description of the nuclear profiles as compared to the traditional uniform density or a Woods-Saxon description of the nucleus [70]. In this study, we use the sum of two Gaussian functions to simulate 12C and 16O nuclear profiles. The SOG nuclear charge density () as a function of radial distance () is expanded as,
@@
II.2.2 –clustered nuclear geometry
The 4He nucleus consisting of a pair of protons and a pair of neutrons is known as the –particle. As discussed earlier, nucleons of certain light nuclei can cluster into groups of –particles forming a stable geometrical shape. For this, the nuclei should have ( is a positive integer) number of nucleons. In 12C, three such –particles cluster into forming an equilateral triangle, whereas in 16O, four –particles cluster into a regular tetrahedral arrangement. This –cluster geometry is believed to account for the additional stability of the nucleus. In our study, for the first time, we have implemented such –cluster geometry for 12C and 16O nuclei to study the initial-state effects in particle production through p–C and p–O collisions at the LHC. The technical details of the implementation of –cluster structure for 12C and 16O are discussed below.
@@
For the nucleus, four –particles are situated at the four vertices of a regular tetrahedron with side length 3.42 fm, which makes the rms radius of to be 2.699 fm [39, 16, 38]. One can visualize the nucleons inside the 16O nuclei as depicted in Fig. 1. Nucleons inside the –particle are sampled using the following Woods-Saxon density profile in terms of a three-parameter Fermi (3pF) distribution,
(2) Here, is the nuclear charge density at a radial distance from the center of the nucleus. refers to the mean radius of the nucleus, is the nuclear deformation parameter, and is the skin depth. The values of these parameters are chosen for the 4He nucleus as = 0.964 fm, = 0.517, and = 0.322 fm, which corresponds to the rms radius of 1.676 fm for the –particle. To include a finite volume effect for each nucleon, the minimum separation distance between any two nucleons is set to be 0.4 fm. The spatial orientation of the tetrahedron is randomized before each collision for both the target and projectile nuclei.
@@
For the 12C nucleus, three –particles arrange themselves at the corners of an equilateral triangle of side length 3.10 fm [49, 72]. One can visualize the nucleons inside the 12C nuclei as depicted in Fig. 1. These parameters lead to a rms radius of 2.47 fm for the 12C nucleus. The nucleons inside the –particles are sampled using a similar procedure as discussed above.

### v2

@@ Abstract
Long-range multi-particle correlations in heavy-ion collisions have shown conclusive evidence of the hydrodynamic behavior of strongly interacting matter and are associated with the final-state azimuthal momentum anisotropy. In small collision systems, azimuthal anisotropy can be influenced by the hadronization mechanism and residual jet-like correlations. Thus, one of the motives of the planned p–O and O–O collisions at the LHC and RHIC is to understand the origin of small system collectivity. As the anisotropic flow coefficients () are sensitive to the initial-state effects including nuclear shape, deformation, and charge density profiles, studies involving 12C and 16O nuclei are transpiring due to the presence of exotic (4He) clusters in such nuclei. In this study, for the first time, we investigate the effects of nuclear –clusters on the azimuthal anisotropy of the final-state hadrons in p–C and p–O collisions at TeV within a multi-phase transport model framework. We report the transverse momentum () and pseudorapidity () spectra, participant eccentricity () and triangularity (), and estimate the elliptic flow () and triangular flow () of the final-state hadrons using the two-particle cumulant method. These results are compared with a model-independent Sum of Gaussians (SOG) type nuclear density profile for 12C and 16O nuclei.
@@ I Introduction
The world’s most powerful particle accelerators, the Relativistic Heavy-Ion Collider (RHIC) at BNL, USA, and the Large Hadron Collider (LHC) at CERN, Switzerland, offer unparalleled experimental capabilities to perform high-energy hadronic and nuclear collisions to study the behavior of nuclear matter under extreme conditions of temperature and energy density. Such ultra-relativistic heavy-ion collisions produce a state of locally thermalized and deconfined partonic medium, known as the quark-gluon plasma (QGP). While the existence of QGP has long been established in heavy-ion collisions, its presence in small collision systems is a matter of intense research, recently. Thanks to the recent measurements of heavy-ion-like features in high multiplicity pp collisions such as strangeness enhancement [1], ridge-like structure [3, 2], and radial flow effects [4, 5, 6], the possible existence of QGP droplets in small collision systems needs to be re-examined. As p–C and p–O collisions fill the multiplicity gap between pp to p–Pb and peripheral Pb–Pb collisions, they provide a perfect system size for studying several heavy-ion-like effects in small collision systems at the RHIC and LHC energies [7, 8].
@@
The anisotropic flow coefficients, such as elliptic flow () and triangular flow (), are found to be sensitive to the initial nuclear distribution, nuclear shape deformation, and fluctuations in the nuclear overlap region [16, 14, 15]. In Xe–Xe collisions at the LHC, the presence of a deformed nuclear structure gives rise to a higher value of elliptic flow in the most central collisions compared to Pb–Pb collisions [17, 18, 20, 19]. In addition, a larger density fluctuation due to a smaller number of participating nucleons in Xe–Xe collisions gives rise to a larger value of triangular flow in contrast to Pb–Pb collisions [17, 18]. A similar study using a multi-phase transport model (AMPT) in U–U collisions at the RHIC energy shows that the values of anisotropic flow coefficients change with a change in the deformation parameter of the nuclear distribution [14]. However, the impact of this deformation is larger for elliptic flow in contrast to the triangular flow [14, 15, 21]. Additionally, recent observations of ridge-like structures in p–Pb collisions have brought significant attention to asymmetric collisions, which are customarily considered to provide a baseline measurement for the cold nuclear effects [22, 23, 24, 25, 6], where the presence of strong long-range multi-particle correlations are not expected. Further, the measured values of anisotropic flow coefficients in p–Pb collisions are comparable to that of peripheral Pb–Pb collisions [27, 26, 28]. Although, in p–Pb collisions, the ordering of the magnitude of flow coefficients, i.e., , is not as strong as in Pb–Pb collisions, the bare presence of signatures of collectivity in p–Pb collisions makes the study of asymmetric collisions a vital contribution of the heavy-ion physics community.
@@
With all these tempting motivations, in recent years, several studies have been performed involving collisions of 12C, 16O and 20Ne nuclei [40, 41, 39, 35, 36, 37, 16, 38, 42, 43, 44]. It includes the studies based on different hydrodynamic models [45, 46, 47, 42], Glauber Monte Carlo [35, 36, 37], parton energy loss [48], and jet quenching effects [49]. Some of these studies also investigate the possible signatures of –clusters in 12C and 16O nuclei by studying the final-state particle production and anisotropic flow measurements [40, 41, 39, 51, 42, 52, 53, 54, 16, 50]. In Ref. [40], an enhanced value of triangular flow in C–Au collisions is observed for an –clustered 12C nucleus. In Ref. [16], authors show that the presence of an –clustered structure in 16O nucleus leads to a high value of triangular flow in most central O–O collisions, which are well observed in the ratio . In addition, an away-side broadening in the two-particle correlation function is also reported for O–O collisions with –clusters as compared to the Woods-Saxon nuclear profile [16]. Although there have been a few studies that aim to establish a clear signature of the presence of –clusters in 16O and 12C nucleus, the choice of p–O and p–C is novel to this study. In addition, the studies of anisotropic flow in asymmetric collision systems like p–O and p–C can accord the observations of collectivity in p–Pb collisions, which can be an interesting and a new addition to the existing results of the geometry scan program at RHIC [56, 55, 57, 58].
@@
In this study, we incorporate an –cluster-type initial geometry for 12C and 16O nuclei, and report the first measurements of transverse momentum () and pseudorapidity () spectra, participant eccentricity () and triangularity (), and elliptic flow () and triangular flow () of the final-state hadrons, and their scaling in p–O and p–C collisions at TeV using AMPT. We use the two-particle cumulant method with a relative pseudorapidity cut to minimize the effects of non-flow correlations. These results are compared with a model-independent Sum of Gaussians (SOG) type nuclear density profile for 12C and 16O nuclei as an unclustered description of nuclei is essential for a good understanding of the effects in the final state.
@@
The paper is organized as follows. We start with a brief introduction and motivation of the study in Section I followed by a discussion on the implementation of various nuclear density profiles in Section II. In Section III, we present and discuss the results. Finally, in Section IV, we summarize the study with important findings and present a brief outlook for the future. A detailed description of the event generation using AMPT and the anisotropic flow estimation using the two-particle Q-cumulant method are discussed in Appendix A.
II NUCLEAR DENSITY PROFILE IMPLEMENTATION
@@
This study implements two different nuclear density profiles for the 12C and 16O nuclei: –cluster type geometrical distribution and a model-independent Sum of Gaussians type nuclear density profile. The technical details regarding the implementation of the two density profiles are discussed below.
II.0.1 Sum Of Gaussians (SOG) density profile
@@
To represent the nuclear charge density and extract charge density parameters in a model-independent fashion as approximated in experiments, one fits the nuclear densities with a sum of Gaussian functions (SOG) [59, 60]. SOG can accurately represent complex nuclear densities by exploiting the parameters of individual Gaussian functions, which makes SOG applicable to a broad list of nuclei. SOG can produce smooth and continuous form of the nuclear density profiles, consequently simplifying many theoretical calculations due to the well-known properties of the Gaussian distribution function. In addition, SOG can produce a more realistic and precise description of the nuclear profiles as compared to the traditional uniform density or a Woods-Saxon description of the nucleus [59]. In this study, we use the sum of two Gaussian functions to simulate 12C and 16O nuclear profiles. The SOG nuclear charge density () as a function of radial distance () is expressed as,
@@
II.0.2 –clustered nuclear geometry
The 4He nucleus consisting of a pair of protons and a pair of neutrons is known as the –particle. As discussed earlier, nucleons of certain light nuclei can cluster into groups of –particles forming a stable geometrical shape. For this, the nuclei should have ( is a positive integer) number of nucleons. In 12C, three such –particles cluster into forming an equilateral triangle, whereas in 16O, four –particles cluster into a regular tetrahedral arrangement. This –cluster geometry is believed to account for the additional stability of the nucleus. In our study, for the first time, we have implemented such –cluster geometry for 12C and 16O nuclei to study the initial-state effects in particle production through p–C and p–O collisions at the LHC. The technical details of the implementation of –cluster structure for 16O and 12C nuclei are discussed below.
@@
For the nucleus, four –particles are situated at the four vertices of a regular tetrahedron with side length 3.42 fm, which makes the root mean squared (rms) radius of to be 2.699 fm [39, 16, 38]. One can visualize the nucleons inside the 16O nuclei as depicted in Fig. 1. Nucleons inside the –particle are sampled using the following Woods-Saxon density profile in terms of a three-parameter Fermi (3pF) distribution,
(2) Here, is the nuclear charge density at a radial distance from the center of the nucleus. refers to the mean radius of the nucleus, is the nuclear deformation parameter, and is the skin depth. The values of these parameters are chosen for the 4He nucleus as = 0.964 fm, = 0.517, and = 0.322 fm, which corresponds to the rms radius of 1.676 fm for the –particle. To include a finite volume effect for each nucleon, the minimum separation distance between any two nucleons is set to be 0.4 fm [61]. The spatial orientation of the tetrahedron is randomized before each collision for both the target and projectile nuclei.
@@
Here, three –particles arrange themselves at the corners of an equilateral triangle of side length 3.10 fm leading to an rms radius of 2.47 fm for the 12C nucleus [52, 62]. A depiction of the nucleons inside an -clustered 12C nuclei can be seen in Fig. 1. The nucleons inside the –particles are sampled using a similar procedure as discussed above for the case of nucleus.